OpenStudy (anonymous):

Need some help 25w^2 -25w=-3

OpenStudy (agreene):

\[25w^2-25w=-3\] \[w^2-w =-\frac{3}{25}\] \[w^2-w+\frac{1}{4}=\frac{13}{100}\] \[(w-\frac12)^2=\frac{13}{100}\] sqrt and solve for w, you should find: \[w=\frac12\pm\frac{\sqrt{13}}{10}\]

OpenStudy (anonymous):

Ok I solved using quadratic equation.. I got the answer \[w=\frac{ 1\pm \sqrt{13} }{ 5 } \] Did I do something wrong?

OpenStudy (agreene):

I don't use the quadratic, however our answer are very similar--but i would make sure you did your division and factoring accurately as it seems like we are a bit off from one another

OpenStudy (anonymous):

Alright I will work on it again.

OpenStudy (anonymous):

solved for w

OpenStudy (anonymous):

I have check my work again and got the final answer (I made 1 mistake)\[w=\frac{ 1\pm \sqrt{13}}{ ? }\]

OpenStudy (anonymous):

over 2 i mean not ?

OpenStudy (anonymous):

@SpecialMathStudent Couldn'y you simplify the 5/10 by making 1/2? Because that is what I did.

OpenStudy (anonymous):

Yeah most likely, if its asking you to reduce it to the lowest form.

OpenStudy (mertsj):

\[w=\frac{25\pm \sqrt{625-300}}{50}=\frac{25\pm \sqrt{325}}{50}=\frac{25\pm5\sqrt{13}}{50}=\]

OpenStudy (mertsj):

\[\frac{5\pm \sqrt{13}}{10}\]

OpenStudy (anonymous):

@Mertsj So reducing that to \[\frac{1\pm \sqrt{13}}{2}\] is the wrong answer?

OpenStudy (mertsj):

Or you could split it up into 2 fractions and write: \[\frac{1}{2}\pm\frac{\sqrt{13}}{10}\]

OpenStudy (anonymous):

Id say most likely, stick with the 5 and 10 that should be your answer.

OpenStudy (mertsj):

Yes. That is wrong.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Yeap!

OpenStudy (mertsj):

yw